# Model Breadcrumbs: Scaling Multi-Task Model Merging with Sparse Masks

### v1

@@ 4.3 Validation-Free Setting
Figure 4: The 200-task sequence originates from the ImageNet dataset [12], created by dividing the data into 200 5-class classification tasks. After encountering 10 tasks using the ViT-L-14 model, the best hyperparameters for each method (Breadcrumbs with 85% sparsity and Task Vectors [20]) are selected and fixed. Each point on the plot represents the evaluation of the method over all tasks observed up to that point. With an increasing number of tasks, Model Breadcrumbs consistently outperforms Task Vectors [20] by a substantial margin, highlighting the robustness of hyperparameters in the Model Breadcrumbs approach.
@@
Figure 5: Comparative performance analysis of Model Breadcrumbs and Task Vector [20] methods across varying CLIP model scales (ViT-B-32, ViT-B-16, and ViT-L-14) as the number of tasks increases. The solid line represents the averaged normalized accuracy across all evaluation points. Each data point corresponds to an experiment involving a subset of the 8 tasks under study. Our findings highlight the potential of larger-scale models to mitigate performance degradation and, as seen in Figure 4(b), the capability of Model Breadcrumbs to produce multi-task models that surpass individual fine-tuned models for specific tasks.
@@
In Section 4.2, we compared Model Breadcrumbs and Task Vectors [20] under their respective optimal hyperparameters. These hyperparameters were fine-tuned based on model performance on the validation dataset for each subset of tasks following Ilharco et al. [20]. However, as the number of tasks increases, the search for optimal hyperparameters becomes increasingly resource-intensive. Furthermore, the need for a validation set from each task being added can be restrictive due to privacy concerns or due to the unavailability of additional validation data. Thus, we consider a new setting where hyperparamters are tuned based on a few tasks, and subsequent tasks are added using these pre-determined hyperparameters.
@@
The results are shown in Figure 3. Remarkably, our experiments reveal that the hyperparameters of Model Breadcrumbs exhibit a high degree of generalizability. Specifically, for the ViT-B-32 model when considering scenarios involving three tasks and beyond, up to the 8-task scenario, the optimal hyperparameters remain consistent. Moreover, for the ViT-L-14 model, the hyperparameters do not change beyond the 1 task scenario. This remarkable stability underscores the robustness and versatility of Model Breadcrumbs. We observer that on the other hand the approach Task Vectors [20] can quickly collapse in performance.
@@
Motivated by these results, we extended the evaluation to a much longer task sequence using ViT-L-14 model. We split the ImageNet data [12] into 200 tasks, each classifying 5 classes. After finding optimal hyperparameters for both Model Breadcrumbs and Task Vectors using 10 tasks, we kept these hyperparameters and incrementally merged all 200 tasks to create a multi-task model. As seen in Figure 4, the observed trend remains, with Model Breadcrumbs (85% sparsity) consistently outperforming Task Vectors [20] by a significant margin as the number of tasks increases. This showcases the generalizability of the hyperparameters for the Model Breadcrumbs approach.
@@
The practical implication of this stability in hyperparameter settings is that, in practice, we can rely on a relatively small number of tasks to determine optimal hyperparameters when applying Model Breadcrumbs to diverse multi-task learning scenarios. This simplifies the implementation process, reduces the need for extensive hyperparameter tuning, and contributes to the framework’s practicality and ease of use. In contrast, Task Vectors [20] do not exhibit the same level of hyperparameter stability. Consequently, this fundamental divergence between Model Breadcrumbs and Task Vectors [20] underlines the substantial advantage of Model Breadcrumbs in real-world multi-task learning scenarios.
@@ 4.4 Effect of Scale
Figure 6: Comparison of Model Breadcrumbs and Task Vectors [20] in the merger of task pairs, revealing improved accuracy on both tasks and a higher frequency of multi-task models surpassing individual fine-tuned accuracy levels when employing Model Breadcrumbs.
In this section, we explores the impact of using larger CLIP models on our analysis, comparing the performance of ViT-B-32, ViT-B-16, and ViT-L-14 models. For each model type, the optimal Model Breadcrumbs were found at 85% sparsity. As shown in Figure 5, the adoption of larger models significantly improves the performance of both our proposed Model Breadcrumbs method and the Task Vector [20] baseline. Moreover, as more tasks are introduced, the capacity to construct better-performing multi-task models grows, with larger-scale models demonstrating superior results.
@@
Moreover, Figure 4(b) highlights that the performance decline observed when merging either Model Breadcrumbs or Task Vectors [20] can be significantly mitigated by adopting larger-scale models. Notably, for the ViT-L-14 model, merging Model Breadcrumbs for certain tasks can result in multi-task models that either match or surpass the performance of individual fine-tuned models. To delve deeper into this phenomenon, we conducted a closer examination of task merger for ViT-L-14, considering the two tasks scenario.
As we can see in Figure 6, when adding pairs of tasks via Model Breadcrumbs and Task Vectors [20], the merger generally leads to improved performance on both tasks, resulting in a single model that is competitive and often superior to using two specialized fine-tuned models. Furthermore, for the same task pairs, Model Breadcrumbs consistently produces multi-task models that surpass their equivalent Task Vectors [20] versions. Notably, Model Breadcrumbs mergers generate a higher number of multi-task models where both tasks exceeded their respective fine-tuned accuracy levels. This highlights the potential of Model Breadcrumbs not only to maintain but also to enhance task-specific performance within a multi-task framework. We further examine this concept in the next section.
4.5 Task Improvement via Model Merging
@@
Fine-tuned + Task Vectors 88.6 ±$\pm 0.59$ 76.4 ±$\pm 0.40$ 52.1 ±$\pm 0.32$ 93.5 ±$\pm 0.31$
@@
Table 2: Model merging enhances fine-tuned models. Specifically, the merger of Breadcrumbs yields higher-performing models without requiring additional training data or combining with models trained on similar data. Values represent the average performance over 20 runs, followed by the standard error.
Motivated by the insights from Figure 6, we explore the potential of enhancing fine-tuned model performance solely through model merging. We fine-tune the T5-base model [44] for four GLUE tasks [53] (discussed in Section 4.1) based on benchmarks used by [20, 56]. We then merge six publicly available T5-base models (IMDB [34], RACE [26], QASC [23], MultiNews [16], SQuAD [46], and CommonGen [30]) with each of them. Appendix B provides more details on fine-tuned models and the fine-tuning process.
The results, presented in Table 2, evaluates various approaches, including Zeroshot, pure fine-tuning, utilizing Task Vectors [20] derived from newly added tasks in the fine-tuned models, and employing Breadcrumbs of the new tasks added to the fine-tuned models. The findings demonstrate that incorporating Breadcrumbs from these new tasks effectively enhances the performance of our fine-tuned models, surpassing all other considered approaches. Importantly, this improvement is achieved without the need for additional training or requiring data from the exact same dataset. This approach underscores the versatility and effectiveness of utilizing Model Breadcrumbs to improve task performance across diverse tasks.
@@ 4.6 Ablations
Figure 7: Performance comparison of the Model Breadcrumbs against alternative masking choices, reveals: (1) Model Breadcrumbs yields a higher distribution of high-performance multi-task models, underlining its robustness towards hyperparameter perturbations. (2) Model Breadcrumbs produces the highest performing multi-task model. The number on top of each violin indicates the performance of the highest performing model of that setting.
@@
Figure 8: Comparison of Cosine Similarity Between Tasks in Model Breadcrumbs and Task Vectors. The figure illustrates the cosine similarity distribution among tasks, highlighting how Model Breadcrumbs enforces greater orthogonality, leading to reduced interference during model merging.
In this section, we perform ablations to examine alternative design decisions within the Model Breadcrumbs method. Specifically, we explore different approaches for constructing the masking operation, namely: 1. Bottom-Weight Masking: Masking only the bottom-most smallest absolute magnitude weights per layer. 2. Top-Weight Masking: Masking only the top largest absolute magnitude weights per layer.
@@
We compare these alternatives to the full Model Breadcrumbs approach, which encompasses both (1) and (2), as well as the Task Vectors [20] method, which lacks any masking. In our investigation, we conduct a grid search to identify the optimal hyperparameters for each of the four configurations. We assess the resulting multi-task models on 8 tasks discussed in Section 4.1. The results are shown in Figure 7.
@@
Our findings reveal two key insights: [label=()] 1. both forms of weight masking, as employed in Model Breadcrumbs, are essential for achieving competitive performance. Model Breadcrumbs, which combines both bottom and top weight masking, emerges as the most effective approach. 2. The grid search for hyperparameters within the Model Breadcrumbs approach yields a higher distribution of high-performance multi-task models compared to the other three settings. Furthermore, there is much lower variation in the overall performance distribution of the multi-task models produced by the Model Breadcrumbs. These observations underscore the robustness of Model Breadcrumbs to variations in hyperparameter settings, further enhancing its practicality and reliability in real-world applications.
@@
In Figure 8, we examine the cosine similarity between tasks using Model Breadcrumbs and Task Vectors [20]. Most tasks show orthogonality, indicating minimal side effects upon merging. However, upon closer inspection, semantically similar tasks (e.g., MNIST [27], SVHN [37], and GTSRB [19]) exhibit higher cosine similarity, suggesting non-orthogonality. This similarity could introduce interference during merging. In contrast, Model Breadcrumbs pushes all cosine similarity values closer to zero, reinforcing orthogonality. This reduction in interference could explain the enhance performance of the resulting multi-task models when using Breadcrumbs.
@@ 5 Conclusions
Appendix A Vision: Data and Training
We fine-tune CLIP models [43] using the datasets specified in Section 4.1 (more details are given in Tabel 3). Following a procedure akin to Ilharco et al. [21], our fine-tuning comprises 2000 iterations with a batch size of 128, a learning rate set to 1e-5, and a cosine annealing learning rate schedule with 200 warm-up steps. The AdamW optimizer [32] with a weight decay of 0.1 is employed for optimization.
@@ Appendix A Vision: Data and Training
Appendix B NLP: Data and Training

### v2

@@ 4.3 Validation-Free Setting
Figure 4: The 200-task sequence originates from the ImageNet dataset [13], created by dividing the data into 200 5-class classification tasks. After encountering 10 tasks using the ViT-L-14 model, the best hyperparameters for each method (Breadcrumbs with 85% sparsity and Task Arithmetic [21]) are selected and fixed. Each point on the plot represents the evaluation of the method over all tasks observed up to that point. With an increasing number of tasks, Model Breadcrumbs consistently outperforms Task Arithmetic [21] by a substantial margin, highlighting the robustness of hyperparameters in the Model Breadcrumbs approach.
@@
Figure 5: Comparative performance analysis of Model Breadcrumbs and Task Arithmetic [21] methods across varying CLIP model scales (ViT-B-32, ViT-B-16, and ViT-L-14) as the number of tasks increases. The solid line represents the averaged normalized accuracy across all evaluation points. Each data point corresponds to an experiment involving a subset of the 8 tasks under study. Our findings highlight the potential of larger-scale models to mitigate performance degradation and, as seen in Figure 5(b), the capability of Model Breadcrumbs to produce multi-task models that surpass individual fine-tuned models for specific tasks.
@@
In Section 4.2, we compared Model Breadcrumbs and Task Arithmetic [21] under their respective optimal hyperparameters. These hyperparameters were fine-tuned based on model performance on the validation dataset for each subset of tasks following [21]. However, as the number of tasks increases, the search for optimal hyperparameters becomes increasingly resource-intensive. Furthermore, the need for a validation set from each task being added can be restrictive due to privacy concerns or due to the unavailability of additional validation data. Thus, we consider a new setting where hyperparamters are tuned based on a few tasks, and subsequent tasks are added using these pre-determined hyperparameters.
@@
The results are shown in Figure 3. Remarkably, our experiments reveal that the hyperparameters of Model Breadcrumbs exhibit a high degree of generalizability. Specifically, for the ViT-B-32 model when considering scenarios involving three tasks and beyond, up to the 8-task scenario, the optimal hyperparameters remain consistent. Moreover, for the ViT-L-14 model, the hyperparameters do not change beyond the 1 task scenario. This remarkable stability underscores the robustness and versatility of Model Breadcrumbs. We observer that on the other hand the approach Task Arithmetic [21] can quickly collapse in performance.
@@
Motivated by these results, we extended the evaluation to a much longer task sequence using ViT-L-14 model. We split the ImageNet data [13] into 200 tasks, each classifying 5 classes. After finding optimal hyperparameters for both Model Breadcrumbs and Task Arithmetic using 10 tasks, we kept these hyperparameters and incrementally merged all 200 tasks to create a multi-task model. As seen in Figure 4, the observed trend remains, with Model Breadcrumbs (85% sparsity: $\beta=85\%$, $\gamma=99.3\%$) consistently outperforming Task Arithmetic [21] by a significant margin as the number of tasks increases. This showcases the generalizability of the hyperparameters for the Model Breadcrumbs approach.
@@
The practical implication of this stability in hyperparameter settings is that, in practice, we can rely on a relatively small number of tasks to determine optimal hyperparameters when applying Model Breadcrumbs to diverse multi-task learning scenarios. This simplifies the implementation process, reduces the need for extensive hyperparameter tuning, and contributes to the framework’s practicality and ease of use. In contrast, Task Arithmetic [21] do not exhibit the same level of hyperparameter stability. Consequently, this fundamental divergence between Model Breadcrumbs and Task Arithmetic [21] underlines the substantial advantage of Model Breadcrumbs in real-world multi-task learning scenarios.
@@ 4.4 Effect of Scale
Figure 6: Comparison of Model Breadcrumbs and Task Arithmetic [21] in the merger of task pairs, revealing improved accuracy on both tasks and a higher frequency of multi-task models surpassing individual fine-tuned accuracy levels when employing Model Breadcrumbs.
In this section, we explores the impact of using larger CLIP models on our analysis, comparing the performance of ViT-B-32, ViT-B-16, and ViT-L-14 models. For each model type, the optimal Model Breadcrumbs were found at 90% ($\beta=90\%$, $\gamma=99\%$), 90% ($\beta=90\%$, $\gamma=99.2\%$), 85% ($\beta=85\%$, $\gamma=99\%$) sparsity respectively. As shown in Figure 5, the adoption of larger models significantly improves the performance of both our proposed Model Breadcrumbs method and the Task Arithmetic [21] baseline. Moreover, as more tasks are introduced, the capacity to construct better-performing multi-task models grows, with larger-scale models demonstrating superior results.
@@
Moreover, Figure 5(b) highlights that the performance decline observed when merging either Model Breadcrumbs or Task Arithmetic [21] can be significantly mitigated by adopting larger-scale models. Notably, for the ViT-L-14 model, merging Model Breadcrumbs for certain tasks can result in multi-task models that either match or surpass the performance of individual fine-tuned models. To delve deeper into this phenomenon, we conducted a closer examination of task merger for ViT-L-14, considering the two tasks scenario.
As we can see in Figure 6, when adding pairs of tasks via Model Breadcrumbs and Task Arithmetic [21], the merger generally leads to improved performance on both tasks, resulting in a single model that is competitive and often superior to using two specialized fine-tuned models. Furthermore, for the same task pairs, Model Breadcrumbs consistently produces multi-task models that surpass their equivalent Task Arithmetic [21] versions. Notably, Model Breadcrumbs mergers generate a higher number of multi-task models where both tasks exceeded their respective fine-tuned accuracy levels. This highlights the potential of Model Breadcrumbs not only to maintain but also to enhance task-specific performance within a multi-task framework. We further examine this concept in the next section.
4.5 Target Task Improvement via Model Merging
Table 2: Model merging enhances the fine-tuned models. Specifically, the merger of Breadcrumbs yields higher-performing models without requiring additional training data or combining with models trained on similar data. Values represent the average performance over 20 runs, followed by the standard error.
@@ 4.5 Target Task Improvement via Model Merging
Fine-tuned + Task Arithmetic 88.6 ±$\pm 0.59$ 76.4 ±$\pm 0.40$ 52.1 ±$\pm 0.32$ 93.5 ±$\pm 0.31$
@@
Motivated by the insights from Figure 6, we explore the potential of enhancing the performance of a fine-tuned model for a specific target task solely through model merging. We fine-tune the T5-base model [46] for four GLUE tasks [54] (discussed in Section 4.1) based on benchmarks used by [21, 57]. We then merge six publicly available T5-base models (IMDB [35], RACE [27], QASC [24], MultiNews [17], SQuAD [47], and CommonGen [31]) with each of them. Appendix 0.B provides more details on fine-tuned models and the fine-tuning process.
The results, presented in Table 2, evaluates various approaches, including Zeroshot, pure fine-tuning, utilizing Task Arithmetic [21] derived from newly added tasks in the fine-tuned models, and employing Breadcrumbs of the new tasks added to the fine-tuned models. The findings demonstrate that incorporating Breadcrumbs from these new tasks effectively enhances the performance of our fine-tuned models, surpassing all other considered approaches. Importantly, this improvement is achieved without the need for additional training or requiring data from the exact same dataset. This approach underscores the versatility and effectiveness of utilizing Model Breadcrumbs to improve task performance across diverse tasks.
@@ 4.6 Ablations
Figure 7: Performance comparison of the Model Breadcrumbs against alternative masking choices, reveals: Model Breadcrumbs yields a higher distribution of high-performance multi-task models, underlining its robustness towards hyperparameter perturbations. Model Breadcrumbs produces the highest performing multi-task model. The number on top of each violin indicates the performance of the highest performing model of that setting.
@@
Figure 8: Comparison of Cosine Similarity Between Tasks in Model Breadcrumbs and Task Arithmetic. The figure illustrates the cosine similarity distribution among tasks, highlighting how Model Breadcrumbs enforces greater orthogonality, leading to reduced interference during model merging.
@@
In this section, we perform ablations to examine alternative design decisions within the Model Breadcrumbs method. Specifically, we explore different approaches for constructing the masking operation, namely: 1. Bottom-Weight Masking: Masking only the bottom-most smallest absolute magnitude weights per layer. 2. Top-Weight Masking: Masking only the top largest absolute magnitude weights per layer. We compare these alternatives to the full Model Breadcrumbs approach, which encompasses both (1) and (2), as well as the Task Arithmetic [21] method, which lacks any masking. In our investigation, we conduct a grid search to identify the optimal hyperparameters for each of the four configurations. We assess the resulting multi-task models on 8 tasks discussed in Section 4.1. The results are shown in Figure 7.
@@
Our findings reveal two key insights: (i) both forms of weight masking, as employed in Model Breadcrumbs, are essential for achieving competitive performance. Model Breadcrumbs, which combines both bottom and top weight masking, emerges as the most effective approach. (ii) The grid search for hyperparameters within the Model Breadcrumbs approach yields a higher distribution of high-performance multi-task models compared to the other three settings. Furthermore, there is much lower variation in the overall performance distribution of the multi-task models produced by the Model Breadcrumbs. These observations underscore the robustness of Model Breadcrumbs to variations in hyperparameter settings, further enhancing its practicality and reliability in real-world applications.
@@
In Figure 8, we examine the cosine similarity between tasks using Model Breadcrumbs and Task Arithmetic [21]. Most tasks show orthogonality, indicating minimal side effects upon merging. However, upon closer inspection, semantically similar tasks (e.g., MNIST [28], SVHN [38], and GTSRB [20]) exhibit higher cosine similarity, suggesting non-orthogonality. This similarity could introduce interference during merging. In contrast, Model Breadcrumbs pushes all cosine similarity values closer to zero, reinforcing orthogonality. This reduction in interference could explain the enhance performance of the resulting multi-task models when using Breadcrumbs.
In Figure 9, we demonstrate the impact of hyperparameters on the performance of models using the ViT-B-32 model, assessed across eight vision tasks outlined in Section 4.1. For ease of readability, in this part we use $\beta$ and $\gamma$ to represent how much weights have been masked. Figure 9(a) examines the relationship between $\alpha$ and $\beta$, the primary determinants of task vector sparsity. As $\beta$ increases and more weights are masked, large alphas, which amplify the remaining weights’ contributions, become less tolerable, necessitating lower $\alpha$’s as $\beta$ grows. Upon identifying optimal $\alpha$ and $\beta$ values, we investigate gamma. In Figure 9(b), we depict the relationship between $\beta$ and gamma. Regardless of beta’s value, the $\gamma$ that optimizes a combination of $\alpha$ and $\beta$ tends to hover around 1%, with lower betas allowing for higher gammas and vice versa. Across both figures, we consistently observe that numerous combinations of alpha, beta, and $\gamma$ result in high-performing merged models, as previously noted in Figure 7.
(a) $\alpha$ vs. $\beta$: as $\beta$ grows in magnitude lower $\alpha$’s are required.
(b) $\beta$ vs. $\gamma$: lower $\beta$’s have more tolerance on variations of $\gamma$.
Figure 9: Influence of hyperparameters on model performance using the ViT-B-32 model across eight vision tasks. For ease of readability, in this part we use $\beta$ and $\gamma$ to represent how much weights have been masked. It shows the relationships between $\alpha$ and $\beta$, and between $\beta$ and $\gamma$, highlighting the stability of hyperparameters across the possible combinations.
@@ 5 Conclusions
Acknowledgements
We acknowledge funding from the NSERC Discovery Grant RGPIN-2021-04104 and FRQNT New Scholar. This research was enabled in part by compute resources provided by Digital Research Alliance of Canada (the Alliance) and Calcul Québec.
Appendix 0.A Vision: Data and Training
We fine-tune CLIP models [42] using the datasets specified in Section 4.1 (more details are given in Tabel 3). Following a procedure akin to [22], our fine-tuning comprises 2000 iterations with a batch size of 128, a learning rate set to 1e-5, and a cosine annealing learning rate schedule with 200 warm-up steps. The AdamW optimizer [33] with a weight decay of 0.1 is employed for optimization.
@@ Appendix 0.A Vision: Data and Training
Appendix 0.B NLP: Data and Training
@@
Appendix 0.C TIES Merging
A concurrent study by Yadav et al. [60] presents a method named TIES. Like the Task Arithmetic method [21], TIES initially constructs a set of Task Vectors. These vectors undergo a masking process to eliminate interfering weights, identified as a percentage of overall weights with low magnitudes. The remaining unmasked weights undergo a sign alignment operation to determine their polarity. Finally, a scaled sum merges the task vectors with the pre-trained model.
Our approach differs from TIES in two key aspects. Firstly, we apply masking to both very large and small magnitude weights of the task vectors to minimize interference, whereas TIES focuses solely on small magnitude weights. Secondly, our masking strategy employs layer-wise masking, wherein a certain percentage of weights are masked based on their magnitude relative to the weights within that layer, as opposed to overall masking, which ranks all model weights by magnitude and masks the smallest ones. Notably, in the context of task vectors, overall masking typically targets weights in the early layers [36].
In Figure 10, we compare our method to TIES [60], where at each point: (I) We show results for ViT-B-32 model where we found the best hyper-parameters for that specific number of tasks for each method. (II) We show the average normalized accuracy over all subsets of the 8 tasks detailed in Section 4.1, amounting to a total of $256=2^{8}$ combinations. (III) The evaluation is performed over all 8 tasks at each point. As we can see from Figure 10, Model Breadcrumbs merging consistently outperforms the TIES method at each point, with the performance gap widening as more tasks are considered. This highlights the significant practical performance advantages of Model Breadcrumbs on a larger scale.
Figure 10: Comparison of Model Breadcrumbs and TIES merging methods across tasks, illustrating Model Breadcrumbs’ consistent outperformance, with the performance gap widening as tasks increase. The results underscore the superior practical performance gains of Model Breadcrumbs at scale.